# Vulnerability Webs:

### v2

@@ 3.2 Equilibrium
The discrete-time Markov chain described above is irreducible and aperiodic and hence admits a unique limiting distribution $\pi$, which is the stationary distribution of the Markov chain (Norris, 1997). The stationary distribution $\pi$ can be represented in closed form as an exponential-family distribution (Mele, 2017) and is given by
@@
while the normalizing constant $c_{kk}(\bm{x},\,\bm{z};\,\bm{\theta})$ of the within-type-$k$ distribution is defined as
@@
The limiting distribution $\pi$ of the discrete-time Markov chain represents the long-run distribution of the software development process, conditional on the types $\bm{z}$ of packages. The first term in (9) represents the likelihood of dependencies among packages of the same type, while the second term represents the likelihood of dependencies between packages of different types.
The stationary distribution $\pi$ implies that the incentives of developers are captured by a potential function of the form
@@
where $Q_{kk}\left(\bm{g}_{kk};\,\bm{x},\,\bm{z},\,\bm{\theta}\right)$ is defined in (10) while $Q_{kl}\left(\bm{g}_{kl};\,\bm{x},\,\bm{z},\,\bm{\theta}\right)$ is defined by
@@
In this software development process, each developer team optimally responds to the existing software dependency network. The Nash equilibria of the underlying potential game of the software development process are the networks that maximize the potential function $Q$ (Mele, 2017, 2022; Monderer and Shapley, 1996; Chandrasekhar, 2016; Blume et al., 2013). These are networks such that no developer team is willing to form an additional dependency or delete an existing dependency, when the need arises to update package dependencies. Upon inspecting the stationary distribution $\pi$, it is evident that the Nash equilibria have the highest probability of being observed, because the Nash equilibria maximize $Q$ and hence maximize $\pi$.
@@
$\begin{array}[]{llllll}W(\bm{g},\bm{x},\bm{z},\bm{\theta})&\coloneqq&\displaystyle\sum_{i=1}^{N}U_{i}(\bm{g},\bm{x},\bm{z},\bm{\theta})\vspace{.25cm}\\ &=&Q(\bm{g},\bm{x},\bm{z},\bm{\theta})\,+\,\gamma\,\displaystyle\sum_{k=1}^{K}\sum_{i=1}^{N}\,\sum_{j\neq i}^{N}\,\sum_{r\neq i,j}^{N}g_{ij}g_{ri}z_{ik}z_{jk}z_{rk},\end{array}$ (14)
which is not equal to the potential function $Q(\bm{g},\bm{x},\bm{z},\bm{\theta})$ when $\gamma\neq 0$, so the networks that maximize welfare may not maximize the potential function and are hence not Nash equilibria.
@@ 4 Estimation and Empirical Results
We perform the estimation under the assumption that the observed software dependency network is a sample from the stationary distribution $\pi$ of the discrete-time Markov process described in Section 3.2. This implies that we can estimate the structural parameters based on a single observation of the software dependency network. We sketch a scalable two-step estimation algorithm and refer the interested reader to the appendix for more technical details.
4.1 Estimation
We introduce a scalable approach to estimating directed network formation models with observable and unobservable types. While scalable estimation methods are available for other models, including stochastic block models (Vu et al., 2013), models which assume that links are independent conditional on latent structure (Mele et al., 2023; Athreya et al., 2018; Kitamura and Laage, 2024), and models for undirected networks (Dahbura et al., 2021; Babkin et al., 2020; Schweinberger and Stewart, 2020), no scalable method for models of directed networks is implemented in publicly available and user-friendly software.
To appreciate the computational challenges involved in estimating directed network formation models, note that the loglikelihood function of the parameter vectors $\bm{\theta}$ and $\bm{\eta}$, given a single observation of a software dependency network $\bm{g}$, is
@@ 4.1 Estimation
where $\mathcal{Z}$ is the set of all $K^{N}$ possible configurations of types $\bm{z}$,  $p$ is the multinomial probability of $\bm{z}$ according to (1),  and $\pi$ is the stationary distribution of $\bm{g}$ conditional on observables $\bm{x}$, unobservables $\bm{z}$, and the structural parameter vector $\bm{\theta}$. Maximizing $\ell$ with respect to $\bm{\theta}$ and $\bm{\eta}$ is challenging, because it involves a sum over all $K^{N}$ possible assignments of $N$ packages to $K$ types, which is exponential in $N$. Worse, for each of the $K^{N}$ possible assignments of $N$ packages to $K$ types, the stationary distribution $\pi$ is a function of an intractable normalizing constant, which is the sum of between- and within-type sums, each of which is exponential in the number of pairs of packages $\binom{N_{k}}{2}$ of type $k$ ($k=1,\ldots,K$). Therefore, direct computation of the likelihood function is infeasible unless $N$ is small (e.g., $N\leq 10$).
To develop scalable methods for directed network formation models, we follow Babkin et al. (2020) and decouple the estimation of the unobserved types from the structural utility parameters using a scalable two-step algorithm:
  1. Step 1: Estimate the types of packages.
@@
In the first step, we estimate the types of the packages by assuming that there are no externalities ($\gamma=0$) and bounding $\ell$ from below using variational methods, and then using maximization-minorization (MM) algorithm to increase the lower bound on $\ell$. We then assign each package to the type with the highest posterior probability.
We show in Appendix A that the loglikelihood function $\ell$ can be approximated by the following variational lower bound:
$\begin{array}[]{lllll}\ell_{LB}(\bm{g},\bm{x},\bm{\alpha},\bm{\beta},\bm{\eta};\,\bm{\Xi})\coloneqq\displaystyle\sum_{i=1}^{N}\sum_{j=1}^{N}\sum_{k=1}^{K}\sum_{l=1}^{K}\xi_{ik}\,\xi_{jl}\log\,\pi_{ij,kl}(g_{ij},\bm{x})+\displaystyle\sum_{i=1}^{N}\sum_{k=1}^{K}\xi_{ik}\left(\log\eta_{k}-\log\xi_{ik}\right),\end{array}$ (16)
where $\bm{\Xi}\coloneqq(\xi_{ik})\in[0,1]^{N\times K}$ are auxiliary parameters that serve as approximations of the posterior probabilities that packages $i$ are of type $k$,  $\eta_{k}$ is the probability of being of type $k$,  and $\pi_{ij,kl}$ is the conditional choice probability of a dependency of package $i$ on package $j$, provided $i$ is of type $k$ and $j$ is of type $l$. It is worth noting that $\ell_{LB}$ provides a family of lower bounds on the loglikelihood function $\ell$: Each possible value of the auxiliary parameters $\bm{\Xi}$ provides a lower bound on the loglikelihood function $\ell$. To obtain the best bound on the loglikelihood function $\ell$, we maximize $\ell_{LB}$ with respect to $\bm{\Xi}$.
Maximizing $\ell_{LB}$ with respect to $\bm{\Xi}$ can be cumbersome and time-consuming. Worse, maximization algorithms may be trapped in local maxima of $\ell_{LB}$. Therefore, we construct a surrogate function $M$, which is more convenient to maximize than $\ell_{LB}$:
$M\left(\bm{\Xi};\,\bm{g},\bm{x},\bm{\alpha},\bm{\beta},\bm{\eta},\bm{\Xi}^{(t)}\right)$$\coloneqq$ $\sum_{i=1}^{N}\sum_{j\neq i}^{N}\sum_{k=1}^{K}\sum_{l=1}^{K}\left(\xi_{ik}^{2}\frac{\xi_{jl}^{(t)}}{2\xi_{ik}^{(t)}}+\xi_{jl}^{2}\frac{\xi_{ik}^{(t)}}{2\xi_{jl}^{(t)}}\right)\log\pi_{ij,kl}^{(t)}(g_{ij},\bm{x})$
@@
where $\xi_{ij}^{(t)}$ denotes the estimate of $\xi_{ij}$ at iteration $t$. Relying on the inequality of arithmetic and geometric means and convexity of the negative logarithm, one can show that (4.1) minorizes (16) in the point $\bm{\Xi}^{(t)}$ (see Vu et al., 2013 for further details). Therefore, maximizing (4.1) with respect to $\bm{\Xi}$ forces $\ell_{LB}$ uphill. This step amounts to a quadratic programming problem under box constraints (Stefanov, 2004). This approach along with fast sparse matrix operations, speeds up computations compared with naive variational algorithms that attempt to maximize $\ell_{LB}$ as a function of $\bm{\Xi}$ directly. [^23] See Appendix A for technical details and explanations on how these approximations improve speed of computation. Using these approximations, we can estimate unobserved types at scale while preserving the correlation structure of software dependencies arising from externalities.
@@
In the second step, we estimate the structural parameter vector $\bm{\theta}$ by maximizing the pseudolikelihood, which is the product of conditional choice probabilities of links conditional on the types estimated in the first step (Babkin et al., 2020; Martínez Dahbura et al., 2023).
The two-step algorithm sketched above is scalable, in the sense that it scales from small to large networks, and has been tested on undirected networks with a quarter of million units (Dahbura et al., 2021). We provide additional theoretical and computational details in Appendix A and the Online Appendix.
@@ 4.2 Results
$\log(n)\times$ Edges $\left(\alpha_{1}^{\ell}\right)$ $-0.533$
Externality ($\gamma$) $-0.095$
@@
$\log(n)\times$ Maturity $\left(\beta_{1}^{\ell}\right)$ $-0.123$
Popularity ($\beta_{2}$) 1.944 0.118
@@
$\log(n)\times$Popularity $\left(\beta_{2}^{\ell}\right)$ $-0.18$
Size ($\beta_{3}$) 1.114 $-0.07$
@@
$\log(n)\times$Size$\left(\beta_{3}^{\ell}\right)$ $-0.122$
@@
The results of our estimation are reported in Table LABEL:tbl:results_python. The first column shows estimates of the structural parameters for links of the same (unobservable) type, while the remaining column provides the estimates for links of different types.
We estimate the externality ($\gamma$) to be negative, providing empirical evidence of the interdependence of links suggested by our strategic model. Essentially, our results provide evidence that when developers form connections, they consider the indirect impacts of their actions, at least partially. This behavior leads to a network architecture that might not be as efficient as one where a central planner aims to maximize the overall utility of all participants. Indeed, the externality encourages developers to establish fewer connections than would be ideal from an aggregate utility-maximization perspective. Maintainers see the number of dependencies of a package they consider linking to as a cost, thus leading to linking decisions that decrease the density of the aggregate network. The negative externality has further implications for contagion, as we explain in the next section.
@@
By contrast, the parameters $\beta_{1},\beta_{2},\beta_{3}$ are all positive within types, indicating that developers are more likely to link to packages of the same type of similar maturity, popularity and size. On the other hand, this homophily seems attenuated by the size of the type, as shown by the negative coefficients of the parameters $\beta_{1}^{\ell},\beta_{2}^{\ell},\beta_{3}^{\ell}$.
When forming dependencies with packages of a different type, the effect of the observables is negative for maturity and popularity, while positive for size. We interpret this as evidence of complementarities across types.
One interpretation consistent with these results is that developers use popular features of same-type packages to include in their own code but use infrastructure-style code from different-type packages, which is not necessarily very popular.
@@ 4.3 Model Fit
In Figure 2 we show a boxplot with summaries of the networks obtained from 100 simulations of the model. The summaries show in Figure 2 are natural summaries, because they are the sufficient statistics of the model conditional on the types of packages.  [^25] We start the simulations at the observed network, and we use a Metropolis-Hastings sampler to update the network. We first run a burn-in of 500,000 iterations, and then sample a network every 50,000 iterations. These summaries are normalized, so that a value of zero indicates a perfect match. The red line is the observed network, so a perfect fit corresponds to all the network statistics concentrated around the red line. The boxplots show that the model is able to capture most of the aggregate structural properties of the Python dependency network: most of the simulated statistics are centered around the observed network. The only exception is the externality (i.e. the number of two-paths in our model), however the observed network falls in the $95\%$ confidence interval of the simulation output. We conclude that the estimated model does a good job in reproducing important features of the network.
@@
In the Online Appendix C, we compare the fit of our model to alternative approaches. An important question is whether the unobserved heterogeneity included in our model is needed. To answer this question we compare our estimated model to a model without unobserved types ($K=1$). This corresponds to a standard Exponential Random Graph model (ERGM), as studied in Mele (2017). The ERGM is not able to match the aggregate density of the network, while our model’s simulations are concentrated around the observed number of links. The second question is whether the externalities are important in explaining the topology of the equilibrium network. Therefore, we compare our model to a Stochastic blockmodel with covariates and $K=10$ types. This corresponds to our model with $\gamma=0$. Our simulations show that the our model does a better job matching the externality within types, and should therefore be preferred. Finally, while there is no currently established method to choose the optimal number of types for this class of complex models, we have done a significant effort in setting an appropriate $K$ for estimation. Our practical approach consists of setting the number of types by checking the model’s fit for different values of $K$. In appendix C we show the fit with $K=5$ and $K=15$. These alternative models overestimate or underestimate the externality level, so our model with $K=10$ is preferred.

### v3

@@ 3.2 Equilibrium
The discrete-time Markov chain described above is an irreducible and aperiodic discrete-time Markov chain with finite state space and hence admits a unique limiting distribution $\pi$, which is the stationary distribution of the Markov chain (Norris, 1997). The stationary distribution $\pi$ can be represented in closed form as an exponential-family distribution (Mele, 2017), that is, an exponential-family random graph model (Lusher et al., 2013) with local dependence (Schweinberger and Handcock, 2015), of the form
@@
while the normalizing constant $c_{kk}$ of the within-type-$k$ distribution is
@@
The limiting and stationary distribution $\pi$ of the Markov chain represents the long-run distribution of the software development process, conditional on the types $\bm{z}$ of repositories. The first term in (10) represents the likelihood of dependencies among repositories of the same type, whereas the second term represents the likelihood of dependencies between repositories of different types. The stationary distribution $\pi$ implies that the incentives of developers are captured by a potential function of the form
@@
where $Q_{kk}$ is defined in (11) while $Q_{kl}$ is defined by
@@
In this software development process, each developer team optimally responds to the existing software dependency network. The Nash equilibria of the underlying potential game of the software development process are the networks that maximize the potential function $Q$ (Mele, 2017, 2022; Monderer and Shapley, 1996; Chandrasekhar, 2016; Blume et al., 2013). These are networks such that no developer team is willing to form an additional dependency or delete an existing dependency, when the need arises to update repository dependencies. Upon inspecting the stationary distribution $\pi$, it is evident that the Nash equilibria have the highest probability of being observed, because the Nash equilibria maximize $Q_{kl}$ and so maximize $\pi$.
@@
$\begin{array}[]{llllll}W(\bm{g},\bm{x},\bm{z},\bm{\theta})&\coloneqq&\displaystyle\sum_{i=1}^{N}U_{i}(\bm{g},\bm{x},\bm{z},\bm{\theta})&=&Q(\bm{g},\bm{x},\bm{z},\bm{\theta})\,+\,\gamma\,\displaystyle\sum_{k=1}^{K}\sum_{i=1}^{N}\,\sum_{j\neq i}^{N}\,\sum_{r\neq i,j}^{N}g_{ij}g_{ri}z_{ik}z_{jk}z_{rk}.\end{array}$ (15)
The welfare function is not equal to the potential function $Q$ when $\gamma\neq 0$, so the networks that maximize welfare may not maximize the potential function and are hence not Nash equilibria.
@@ 4 Estimation and Empirical Results
We infer the software development process by assuming that the observed software dependency network is a sample from the stationary distribution $\pi$ of the discrete-time Markov chain described in Section 3.2. We develop a scalable two-step algorithm by decoupling the estimation of the unobserved types from the structural parameter vector $\bm{\theta}$, which we describe Section 4.1. Theoretical guarantees for type recovery and parameter recovery for each step are available: e.g., theoretical guaranteees for type recovery can be found in Schweinberger et al. (2020), whereas theoretical guarantees for parameter recovery given types include consistency (Schweinberger and Stewart, 2020) and asymptotic normality (Stewart, 2025) of maximum likelihood estimators, and rates of convergence for maximum likelihood and pseudo-likelihood estimators (Stewart and Schweinberger, 2025). Empirical results are presented in Section 4.2.
4.1 Two-Step Algorithm
We describe a scalable two-step algorithm for estimating directed network formation models with observable and unobservable types. Note that scalable methods are available for undirected network formation models (Babkin et al., 2020; Dahbura et al., 2021), but no scalable methods are available for directed network formation models.
The likelihood function of the parameter vectors $\bm{\theta}$ and $\bm{\eta}$ of the software development process based on a single observation of a software dependency network $\bm{g}$ is
@@ 4.1 Two-Step Algorithm
where $\mathcal{Z}$ is the set of all $K^{N}$ possible configurations of types $\bm{z}$,  $p$ is the multinomial probability of $\bm{z}$ according to (1),  and $\pi$ is the stationary distribution of $\bm{g}$ conditional on observables $\bm{x}$, unobservables $\bm{z}$, and the structural parameter vector $\bm{\theta}$. Maximizing $\ell$ with respect to $\bm{\theta}$ and $\bm{\eta}$ is challenging, because it involves a sum over all $K^{N}$ possible assignments of $N$ repositories to $K$ types, which is exponential in $N$. Worse, for each of the $K^{N}$ possible assignments of $N$ repositories to $K$ types, the stationary distribution $\pi$ is a function of an intractable normalizing constant, which is the sum of between- and within-type sums, each of which is exponential in the number of ordered pairs of repositories $N_{k}\,(N_{k}-1)$ of type $k$ ($k=1,\ldots,K$). Therefore, direct computation of the likelihood function is infeasible unless $N$ is small (e.g., $N\leq 10$).
To develop scalable approximations of likelihood-based inference, we follow Babkin et al. (2020) and decouple the estimation of the unobserved types from the estimation of the structural parameter vector $\bm{\theta}$ using the following two-step algorithm:
  1. Step 1: Estimate the types of repositories.
@@
We estimate the unobserved types of repositories as follows.
First, we approximate the loglikelihood function $\ell$ by assuming that there are no externalities ($\gamma=0$), which reduces the loglikelihood function of the directed network formation model to the loglikelihood function of a directed stochastic block model and allows us to tap into the large toolbox developed for stochastic block models.
Second, we bound the loglikelihood function $\ell$ with $\gamma=0$ from below by using Jensen’s inequality. Adapting the arguments of Vu et al. (2013) and Babkin et al. (2020) from undirected to directed network formation models reveals that the resulting lower bound $\ell_{LB}$ of the likelihood function $\ell$ is
$\begin{array}[]{lllll}\ell_{LB}(\bm{\theta},\bm{\eta};\,\bm{\Xi})&\coloneqq&\displaystyle\sum_{i=1}^{N}\sum_{j=1}^{N}\sum_{k=1}^{K}\sum_{l=1}^{K}\xi_{ik}\,\xi_{jl}\log\,\pi_{ij,kl}(g_{ij},\bm{x}_{ij})+\displaystyle\sum_{i=1}^{N}\sum_{k=1}^{K}\xi_{ik}\left(\log\eta_{k}-\log\xi_{ik}\right),\end{array}$ (17)
where $\bm{\Xi}\coloneqq(\xi_{ik})\in[0,1]^{N\times K}$ are auxiliary parameters, $\eta_{k}$ is the prior probability of being of type $k$,  and $\pi_{ij,kl}$ is the conditional choice probability of a dependency of repository $i$ on repository $j$, provided $i$ is of type $k$ and $j$ is of type $l$ with covariates $\bm{x}_{ij}=(\bm{x}_{i},\bm{x}_{j})$. The value $\xi_{ik}$ serves as approximations of the posterior probability of repositories $i$ being of type $k$.
Third, we maximize the lower bound $\ell_{LB}$ with respect to $\bm{\Xi}$ to obtain the best bound on the loglikelihood function $\ell$. Variational algorithms that directly maximize $\ell_{LB}$ with respect to $\bm{\Xi}$ risk being trapped in local maxima and can be time-consuming. To facilitate maximization of $\ell_{LB}$, we construct a surrogate function $M$, which is more convenient to maximize than $\ell_{LB}$. Vu et al. (2013) and Babkin et al. (2020) propose the following surrogate function, which can be obtained using the arithmetic-geometric mean inequality along with concavity properties of logarithms:
$M\left(\bm{\Xi};\,\bm{\theta},\bm{\eta}^{(t)},\bm{\Xi}^{(t)}\right)$$\coloneqq$ $\sum_{i=1}^{N}\sum_{j\neq i}^{N}\sum_{k=1}^{K}\sum_{l=1}^{K}\left(\xi_{ik}^{2}\frac{\xi_{jl}^{(t)}}{2\xi_{ik}^{(t)}}+\xi_{jl}^{2}\frac{\xi_{ik}^{(t)}}{2\xi_{jl}^{(t)}}\right)\log\pi_{ij,kl}^{(t)}(g_{ij},\bm{x}_{ij})$
@@
where $\xi_{ij}^{(t)}$ denotes the estimate of $\xi_{ij}$ at iteration $t$. Maximizing (4.1) with respect to $\bm{\Xi}$ forces $\ell_{LB}$ uphill and can be implemented by quadratic programming using fast sparse matrix operations. The resulting minorization-maximization (MM) algorithm is less prone to being trapped in local maxima than variational algorithm that directly maximize $\ell_{LB}$ with respect to $\bm{\Xi}$.
Finally, we assign repository $i$ to type $\widehat{k}=\mbox{argmax}_{k}\;\widehat{\xi}_{i,k}$, where $\widehat{\xi}_{i,k}$ is an approximation of the posterior probability of $i$ being of type $k$.
@@
We estimate the structural parameter vector $\bm{\theta}$ by maximizing the pseudo-likelihood function of $\bm{\theta}$, which is the product of conditional choice probabilities of links conditional on the types estimated in Step 1 (Babkin et al., 2020; Martínez Dahbura et al., 2023). Maximum pseudo-likelihood estimators are more scalable than maximum likelihood estimators and are therefore preferred on computational grounds.
The resulting two-step algorithm is scalable, in that it can handle small and large networks, and it has been tested on undirected networks with a quarter of million units (Dahbura et al., 2021). Section 4 provides theoretical guarantees for type and parameter recovery in Steps 1 and 2.
@@ 4.2 Results
$\log(n)\times$ Edges ($\alpha_{2})$-0.533
Externality ($\gamma$) --0.095
@@
$\log(n)\times$ Maturity ($\beta_{1}^{\ell}$) --0.123
Popularity ($\beta_{2}$) 1.114 -0.07
@@
$\log(n)\times$Popularity ($\beta_{2}^{\ell}$) --0.122
Size ($\beta_{3}$) 1.944 0.118
@@
$\log(n)\times$Size ($\beta_{3}^{\ell}$) --0.18
@@
The results of our estimation are reported in Table 5. The first column shows estimates of the structural parameters for links of the same (unobservable) type, while the remaining column provides the estimates for links of different types.
We estimate the externality ($\gamma$) to be negative, providing empirical evidence of the interdependence of links suggested by our strategic model. Essentially, our results provide evidence that when developers form connections, they consider the indirect impacts of their actions, at least partially. This behavior leads to a network architecture that might not be as efficient as one where a central planner aims to maximize the overall utility of all participants. Indeed, the externality encourages developers to establish fewer connections than would be ideal from an aggregate utility-maximization perspective. Maintainers see the number of dependencies of a repository they consider linking to as a cost, thus leading to linking decisions that decrease the density of the aggregate network. The negative externality has further implications for contagion, as we explain in the next section.
@@
By contrast, the parameters $\beta_{1},\beta_{2},\beta_{3}$ are all positive within types, indicating that developers are more likely to link to repositories of the same type of similar maturity, popularity and size. On the other hand, this homophily seems attenuated by the size of the type, as shown by the negative coefficients of the parameters $\beta_{1}^{\ell},\beta_{2}^{\ell},\beta_{3}^{\ell}$.
When forming dependencies with repositories of a different type, the effect of the observables is negative for maturity and popularity, while positive for size. We interpret this as evidence of complementarities across types.
One interpretation consistent with these results is that developers use popular features of same-type repositories to include in their own code but use infrastructure-style code from different-type repositories, which is not necessarily very popular.
@@ 4.3 Model Fit
In Figure 2 we show a boxplot with summaries of the networks obtained from 100 simulations of the model. The summaries show in Figure 2 are natural summaries, because they are the sufficient statistics of the model conditional on the types of repositories.  [^24] We start the simulations at the observed network, and we use a Metropolis-Hastings sampler to update the network. We first run a burn-in of 500,000 iterations, and then sample a network every 50,000 iterations. These summaries are normalized, so that a value of zero indicates a perfect match. The red line is the observed network, so a perfect fit corresponds to all the network statistics concentrated around the red line. The boxplots show that the model is able to capture most of the aggregate structural properties of the Python dependency network: most of the simulated statistics are centered around the observed network. The only exception is the externality (i.e. the number of two-paths in our model), however the observed network falls in the $95\%$ confidence interval of the simulation output. We conclude that the estimated model does a good job in reproducing important features of the network.
@@
In the Online Appendix B, we compare the fit of our model to alternative approaches. An important question is whether the unobserved heterogeneity included in our model is needed. To answer this question we compare our estimated model to a model without unobserved types ($K=1$). This corresponds to a standard Exponential Random Graph model (ERGM), as studied in Mele (2017). The ERGM is not able to match the aggregate density of the network, while our model’s simulations are concentrated around the observed number of links. The second question is whether the externalities are important in explaining the topology of the equilibrium network. Therefore, we compare our model to a Stochastic blockmodel with covariates and $K=10$ types. This corresponds to our model with $\gamma=0$. Our simulations show that the our model does a better job matching the externality within types, and should therefore be preferred. Finally, while there is no currently established method to choose the optimal number of types for this class of complex models, we have done a significant effort in setting an appropriate $K$ for estimation. Our practical approach consists of setting the number of types by checking the model’s fit for different values of $K$. In appendix B we show the fit with $K=5$ and $K=15$. These alternative models overestimate or underestimate the externality level, so our model with $K=10$ is preferred.